### v2

# Local quantum channels giving rise to quasi-local Gibbs states

Abstract
We study the steady-state properties of quantum channels with local Kraus operators. We consider a large family that consists of general ergodic 1-local (non-interacting) terms and general 2-local (interacting) terms. Physically, a repeated application of these channels can be seen as a simple model for the thermalization process of a many-body system. We study its steady state perturbatively, by interpolating between the 1-local and 2-local channels with a perturbation parameter $\epsilon$. We prove that under very general conditions, these states are Gibbs states of a quasi-local Hamiltonian. Expanding this Hamiltonian as a series in $\epsilon$, we show that the $k$-th order term corresponds to a $(k+1)$-local interaction term in the Hamiltonian, which follows the same interaction graph as the channel. We also prove a complementary result suggesting the existence of an interaction strength threshold, under which the total weight of the high-order terms in the Hamiltonian decays exponentially fast. For sufficiently small $\epsilon$, this implies both exponential decay of local correlation functions and a classical algorithm with quasi-polynomial runtime for computing expectation values of local observables in such steady states. Finally, we present numerical simulations of various channels that support our theoretical results.
1 Introduction
The study of open quantum systems governed by local dynamics has gained significant attention over the past few decades. Such systems are prevalent in realistic scenarios, as physical systems can never be completely isolated from their surroundings. As a result of their interaction with the environment, they are often subject to non-reversible dynamics, which sends the system to a fixed point [1]. Such states are often referred to as steady states or non-equilibrium steady states (NESS). Like ground states in closed Hamiltonian systems, steady states of many-body open quantum systems exhibit a rich variety of phenomena such as quantum criticality [2, 3, 4], topological phases [5, 6, 7, 8], and chaos [9, 10]. They can also be used as a resource for quantum computation [11, 12, 13], and are crucial for our understanding and modeling of the noise in quantum computers [14, 15, 16].
More recently, many questions and results about open systems, and steady states in particular, originate from a fruitful exchange of ideas with the field of quantum information and quantum computation. A central question in this context, is the preparation of a Gibbs state on a quantum computer, a task also known as Gibbs sampling. This problem has gathered much attention in recent years (see, for example, Refs. [11, 13, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]). In many of these works, one attempts to approach a desired Gibbs state by driving the system to a fixed point of a discrete, stochastic map, which can be implemented on a quantum computer. This naturally leads to the following question:
Under what conditions does the steady state of a discrete, local and stochastic quantum map approximate a Gibbs state of a local Hamiltonian?
Given a local Hamiltonian $H$ on a lattice, its Gibbs state at inverse temperature $\beta=1/k_{B}T$ is given by
$\rho_{H}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{Z}e^{-\beta H},\qquad Z\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathrm{Tr}(e^{-\beta H}).$ (1)
Physically, Gibbs states describe the thermal equilibrium state of a system described by a Hamiltonian $H$, which is in contact with a heat bath of inverse temperature $\beta$. They can also be viewed as quantum generalizations of Boltzmann machines [31, 32], and are therefore important for quantum machine learning. Strictly speaking, any quantum state with a full rank can be written as a Gibbs state $\rho=e^{-H}$ for $H\stackrel{{\scriptstyle\mathrm{def}}}{{=}}-\log(\rho)$. We would therefore like to understand under which conditions does the local structure of the stochastic dynamics imply a local Gibbs Hamiltonian, and more generally, what is the relation between the locality properties of this map and the locality of the Gibbs Hamiltonian. Given the ubiquity of Gibbs states, which under very broad thermodynamical assumptions are the natural equilibrium state of any system, it is reasonable to expect that such local Gibbs Hamiltonians will be the results of many stochastic local maps.
Our question is not only interesting from a theoretical point of view, but also from a practical one. Local stochastic quantum maps can generally be implemented on a quantum computer, and as such, if they generate an interesting Gibbs state, they effectively constitute a quantum algorithm for Gibbs state preparation. In this context, there has been an ongoing effort for constructing such algorithms by designing a local Lindbladian whose steady state is a Gibbs state of a given local Hamiltonian [18, 19, 26, 29, 33, 34]. For example, in Refs. [26, 33], modified versions of the Davis generators [35] introduce such Lindbladians, which consist of terms with a small effective support. It may very well be that a better understanding of the link between the local stochastic maps and the structure of the Gibbs Hamiltonian that they generate will lead to more direct ways of constructing such Gibbs sampling algorithms.
In this work, we use an elementary proof to identify a large family of such maps. We consider $2$-local Kraus maps defined on arbitrary interaction graph and a perturbation parameter $\epsilon\in[0,1)$ such that at $\epsilon=0$, the system is dissipative and non-interacting. We show that under fairly weak assumptions, the steady state of the system is the Gibbs state of a Hamiltonian $H(\epsilon)=\sum_{k\geq 0}\epsilon^{k}H_{k}$, where $H_{k}$ is a sum of terms that are $(k+1)$-geometrically local with respect to the underlying interaction graph. We note that while this result is intuitive, it is not straightforward to prove. A perturbation theory for the steady state equation will yield an expansion of the steady state in powers of $\epsilon$. This expansion will surely contain some fingerprints of the underlying local dynamics; its first few terms are expected to be sums of few operators with a simple tensorial structure. However, the Gibbs Hamiltonian is the logarithm of this operator — a highly non-local mapping — and it is not clear how the local structure of the steady state will descend to the Hamiltonian. In particular, it is entirely not clear how local geometry in $H$ emerges. Our proof, manages to circumvent all this difficulty without delving into an intricate bookkeeping.
If the norm of the $H_{k}$ terms grows at most exponentially in $k$, it follows that for sufficiently small $\epsilon$, the Hamiltonian $H(\epsilon)$ is quasi-local, i.e., the weight of its $k$-local terms decays exponentially. While we cannot prove that this is always the case, we give numerical and analytical evidences to support that. Using numerical simulations of 1D systems, we show that in all the considered cases, $H(\epsilon)$ is always quasi-local, for all $\epsilon\in[0,1)$. In addition, we show the following analytical result. We consider the expectation value of an arbitrary local observable $A$ in the steady state $\rho(\epsilon)$, and expand it perturbatively in $\epsilon$. We obtain a “Heisenberg-like” expression for the expectation value of $A$ in the stead-state in the form of a series of operators $A_{0},A_{1},A_{2},\ldots$ such that ${\langle{A}\rangle}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathrm{Tr}(A\rho(\epsilon))=\sum_{k=0}^{\infty}\epsilon^{k}\mathrm{Tr}(A_{k}\rho_{0})$, where $\rho_{0}$ is the steady state of the $\epsilon=0$ system (i.e., the non-interacting system), and the operators $A_{k}$ satisfy the following properties: 1) $A_{0}=A$ and the support of $A_{k}$ is given by the support of all $k$-steps paths starting from the support of $A$, and 2) the norm of $A_{k}$ grows at most as $O(e^{6k})$. Consequently, for sufficiently small $\epsilon<\epsilon_{0}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}1/e^{6}$, the perturbation series for ${\langle{A}\rangle}$ converges exponentially fast, which is consistent with $H(\epsilon)$ being quasi-local. This result is independent of the underlying interaction graph $G$. If we further assume that the system is defined on a $D$-dimensional lattice, then for $\epsilon<\epsilon_{0}$, it follows that ${\langle{A}\rangle}$ can be approximated up to an additive error of $\delta\cdot{\|A\|}$ using a classical algorithm that runs in time $e^{O(\log^{D}(1/\delta))}$. It is important to note that in the limit $\epsilon\rightarrow 1$, the channel becomes strongly interacting and can efficiently prepare ground states of frustration-free Hamiltonians [11, 12]. Combined with our second result, this suggests a phase transition in the complexity of the channel somewhere in $[\epsilon_{0},1)$. We believe that this transition point might be less universal and depend on the underlying graph and local channels.
The structure of the rest of the paper is as follows. In Sec. 2 we provide some basic theoretical background about quantum maps and Gibbs state, which will allow us to state and prove our main theorem. In Sec. 3, we describe the family of systems that we consider, and give a formal statement of our main results, which are then proved in Sec. 4. In Sec. 5 we present our numerical experiments that support the locality of $H(\epsilon)$, and in Sec. 6 we offer our conclusions and outlook for future research.
2 Background and notation
In this work we consider many-body quantum systems of $n$ qubits that are located on the vertices of an interaction graph $G=(V,E)$ with $n=|V|$. We denote the Hilbert space of the underlying system by $\mathcal{H}=(\mathbb{C}^{2})^{\otimes n}$, and the set of linear operators on $\mathcal{H}$ by $L(\mathcal{H})$.
Given an operator $O\in L(\mathcal{H})$, we shall denote its operator norm by ${\|O\|}_{\infty}$, which is equal to its largest singular value. The operator norm is equal to the Schatten norms ${\|O\|}_{p}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\big{(}\mathrm{Tr}|O|^{p}\big{)}^{1/p}$ with $p\to\infty$ and $|O|\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sqrt{O^{\dagger}\cdot O}$. We shall also make use of the $p=1$ Schatten norm, ${\|O\|}_{1}=\mathrm{Tr}|O|$, known also as the trace norm, and the $p=2$ case ${\|O\|}_{2}=\sqrt{\mathrm{Tr}(O^{\dagger}\cdot O)}$, which is also known as the Frobenius norm or the Hilbert Schmidt norm. We let $\operatorname{supp}(O)$ denote the support of an operator $O$. This is the smallest subset of qubits $S$ such that $O$ can be written as $O=\hat{O}_{S}\otimes\mathbbm{1}_{rest}$ where $\hat{O}_{S}$ defined on the qubits in $S$ and $\mathbbm{1}_{rest}$ is the identity operator on the rest of the qubits. We say that $H\in L(\mathcal{H})$ is a $k$-local Hamiltonian if it can be written as $H=\sum_{\mu}h_{\mu}$, where each $h_{\mu}$ is an Hermitian operator with $\operatorname{supp}(h_{\mu})\leq k$. As Hamiltonians often model physical interactions, we are in particular interested in geometrically $k$-local Hamiltonians, where the interaction terms $h_{\mu}$ have a geometrically local support, i.e., $\operatorname{supp}(h_{\mu})$ is a connected set in $G$. We say that $H=\sum_{\mu}h_{\mu}$ is a geometrically quasi-local Hamiltonian if it can be written as $H=\sum_{k\geq 0}H_{k}$, where $H_{k}$ is geometrically $k$-local Hamiltonian, and ${\|H_{k}\|}_{\infty}$ decays exponentially in $k$.
Given a Hamiltonian $H$, the state $\rho_{H}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{Z}e^{-\beta H}$ with $Z=\mathrm{Tr}(e^{-\beta H})$ is called the Gibbs state of $H$ at inverse temperature $\beta$. This is a physically important state, as it describes the thermal equilibrium state of a system governed by a Hamiltonian $H$, in contact with a heat bath at temperature $T=\frac{1}{k_{B}\beta}$. In this work, we shall always consider Gibbs states of the form $e^{-H}$ by absorbing $\beta$ and $1/Z$ in $H$. Consequently, for any full-rank quantum state $\rho$, we can define $H_{G}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}-\log\rho$ to be its Gibbs Hamiltonian. For a given $\rho$ it is interesting to understand the properties of $H_{G}$, in particular, if it is local or quasi-local with respect to some interaction graph $G$.
A quantum channel is a trace preserving and completely positive (CPTP) linear map over $L(\mathcal{H})$, which we shall usually denote by $\mathcal{E}:L(\mathcal{H})\to L(\mathcal{H})$. Given a channel $\mathcal{E}$, we use the Hilbert-Schmidt inner-product ${\langle{A,B}\rangle}_{HS}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathrm{Tr}(A^{\dagger}B)$ to define its adjoint map $\mathcal{E}^{*}$ by demanding that ${\langle{\mathcal{E}^{*}(A),B}\rangle}_{HS}={\langle{A,\mathcal{E}(B)}\rangle}_{HS}$ for all $A,B\in L(\mathcal{H})$. Note that the trace-preservation property of $\mathcal{E}$ translates into the condition $\mathcal{E}^{*}(\mathbbm{1})=\mathbbm{1}$ for the dual [1].
Of central interest for this work are fixed points of a channel $\mathcal{E}$, which are states that satisfy $\mathcal{E}(\rho)=\rho$. It can be shown that every quantum channel has at least one fixed point that is a proper quantum state (see, for example, Theorem 4.24 in Ref. [36]). We shall refer to these fixed points as steady states.
As in the case of operators, we say that a channel $\mathcal{E}$ (and more generally a super-operator) acts locally on a subset of qubits $S$, if it can be written as =^$\mathcal{E}=\hat{\mathcal{E}}\otimes\mathcal{I}_{rest}$, where $\hat{\mathcal{E}}$ is defined on $L(\mathcal{H}_{S})$ (the space of operators defined on the set of qubits $S$) and $\mathcal{I}_{rest}$ is the identity map on operators that act on the rest of the qubits. We say that a channel $\mathcal{E}$ is $k$-local if it can be written as $\mathcal{E}=\sum_{\mu}\mathcal{F}_{\mu}$, where each $\mathcal{F}_{\mu}$ acts locally on at most $k$ qubits. As a typical example, consider a set of $k$-local quantum gates $\{U_{\mu}\}$ and a corresponding probability distribution $\{p_{\mu}\}$. Then $\mathcal{E}(\rho)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{\mu}p_{\mu}U_{\mu}\rho U_{\mu}^{\dagger}$ is a $k$-local channel.
3 Statement of the main results
Before stating the results formally, we describe the general setup. We consider a many-body system defined on $n$ qubits that sit on the vertices of a graph $G=(V,E)$ with $|V|=n$. Our dynamics is described by a repetitive application of a channel $\mathcal{E}_{\epsilon}$ that depends smoothly on a parameter $\epsilon\in[0,1]$ until we converge to the steady state:
$\rho_{0}\to\mathcal{E}_{\epsilon}(\rho_{0})\to\mathcal{E}^{2}_{\epsilon}(\rho_{0})\to\ldots\to\rho_{\infty}(\epsilon).$
The channel $\mathcal{E}_{\epsilon}$ is defined as follows. We first pick a set of 1-qubit channels $\{\mathcal{D}_{i}\}$, where $\mathcal{D}_{i}$ acts locally on qubit $i$. We further assume that the $\mathcal{D}_{i}$ channels are ergodic [37], i.e., they have a unique, full-rank fixed-point, and no other periodic points (spectrum with modulus one). We denote these fixed points by $\rho_{0}^{(i)}$. Next, we let $\{\mathcal{F}_{e}\}$ be a set of $2$-qubits channels, where $\mathcal{F}_{e}$ acts locally on the pair of qubits that are associated with the edge $e$. We shall refer to the $\mathcal{D}_{i}$ channels as “dissipators” and to $\mathcal{F}_{e}$ as “correlators”.
Given a parameter $\epsilon\in[0,1]$, we combine the dissipators and the correlators to define a 2-qubit channel $\mathcal{E}_{\epsilon}^{(e)}$ that acts locally on the qubits of the edge $e=(i,j)$:
$\mathcal{E}_{\epsilon}^{(e)}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{1}{2}(1-\epsilon)\big{(}\mathcal{D}_{i}+\mathcal{D}_{j}\big{)}+\epsilon\mathcal{F}_{e}$ (2)
For $\epsilon\to 0$, it becomes a sum of two, uncorrelated 1-local channels. For $\epsilon\to 1$, we get $\mathcal{E}_{\epsilon}^{(e)}\to\mathcal{F}_{e}$, which is a general channel on the two qubits. Finally, we pick a probability distribution $\{p_{e}\}$ over the edges of $G$, and define a global channel $\mathcal{E}_{\epsilon}$ by
$\mathcal{E}_{\epsilon}(\rho)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{e\in E}p_{e}\mathcal{E}^{(e)}_{\epsilon}(\rho)=\sum_{e=(i,j)\in E}p_{e}\Big{(}\frac{1}{2}(1-\epsilon)\big{[}\mathcal{D}_{i}(\rho)+\mathcal{D}_{j}(\rho)\big{]}+\epsilon\mathcal{F}_{e}(\rho)\Big{)}.$ (3)
For the sake of concreteness, we give a simple example of the above dynamics, which can be implemented on a quantum computer. Consider a set of $n$ qubits arranged on a circle and choose our dynamics to be translation invariant. We define:
$\mathcal{D}_{i}(\rho)$ =$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathrm{Tr}_{i}[\rho]\otimes\big{(}0.9{|{+}\rangle\langle{+}|}+0.1{|{-}\rangle\langle{-}|}\big{)}_{i},$ $\mathcal{F}_{i,i+1}(\rho)$ =$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\text{C-Z}_{i,i+1}\cdot\rho\cdot\text{C-Z}_{i,i+1},$ (4)
[…]
$\mathcal{E}_{\epsilon}(\rho)=\frac{1-\epsilon}{n}\sum_{i}\mathcal{D}_{i}(\rho)+\frac{\epsilon}{n}\sum_{i}\mathcal{F}_{i,i+1}(\rho).$ (5)
The dynamics described by a repeated application of the channel in Eq. (3) can be thought of simplified model of thermalization for an interacting system locally in contact with a thermal bath. On one hand, the system undergoes local dissipation via the $\mathcal{D}_{i}$ dissipators, and on the other hand, it interacts with itself via the $\mathcal{F}_{e}$ correlators. As the interactions are local, it is expected that the steady state of the system will be described by a Gibbs state of a local or a quasi-local Hamiltonian. Our first result shows that, to a large extent, this is the case.
Theorem 3.1
Consider a many-body quantum system of $n$ qubits as described above, in which the qubits sit on the vertices of the graph $G=(V,E)$, and let $\mathcal{E}_{\epsilon}$ be a local channel as in (3). Then $\mathcal{E}_{\epsilon}$ has a unique, full-rank steady-state $\rho_{\infty}(\epsilon)$ for every $\epsilon\in[0,1)$, which defines a Gibbs Hamiltonian $H_{G}(\epsilon)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}-\log\rho_{\infty}(\epsilon)$. This Hamiltonian is an analytic function of $\epsilon$, which can be written as a series
$H_{G}(\epsilon)=H_{0}+\epsilon H_{1}+\epsilon^{2}H_{2}+\ldots$ (6)
Moreover, for every $k$, the operator $H_{k}$ is a geometrically $(k+1)$-local Hamiltonian with respect to the underlying graph $G$.
From the above theorem, the following corollary easily follows
Corollary 3.2
If there exists a constant $\mu>0$ and a polynomial $C=\operatorname{poly}(n)$ such that ${\|H_{k}\|}_{\infty}\leq C\cdot\mu^{k}$, then for every $\epsilon<\frac{1}{\mu}$, the Gibbs Hamiltonian $H_{G}(\epsilon)$ is quasi-local.
While we suspect that the assumption in the above corollary generally holds, we do not have a proof of it. Nevertheless, we provide two pieces of evidence in its support. Firstly, in Sec. 5, we present a series of numerical experiments of 1D systems, which show that for generic systems, $\epsilon^{k}{\|H_{k}\|}_{\infty}$ decreases exponentially fast in $k$ for every $\epsilon\in[0,1)$. Secondly, we show that an exponential convergence in $k$ for $\epsilon$ below a certain threshold $\epsilon_{0}$ does exist when considering a related quantity. Specifically, we consider the expectation value of an observable $A$ in the steady state:
Theorem 3.3
Consider the system in Theorem 3.1, and assume in addition that the dissipators $\mathcal{D}_{i}$ are of the form $\mathcal{D}_{i}(\rho)=\mathrm{Tr}_{i}(\rho)\otimes W_{i}$, where $W_{i}$ is an arbitrary 1-qubit full-rank state. Let $\rho_{\infty}(\epsilon)=\rho_{0}+\epsilon\rho_{1}+\epsilon^{2}\rho_{2}+\ldots$ be the expansion of the steady state in $\epsilon$ and let $A$ be an observable with $|\operatorname{supp}(A)|=\ell$. Then there exists a sequence of operators $A=A_{0},A_{1},A_{2},\ldots$ such that $\mathrm{Tr}(\rho_{k}A)=\mathrm{Tr}(\rho_{0}A_{k})$ and therefore
${\langle{A}\rangle}(\epsilon)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathrm{Tr}\big{(}\rho_{\infty}(\epsilon)A\big{)}=\sum_{k\geq 0}\epsilon^{k}\mathrm{Tr}(\rho_{0}A_{k}).$ (7)
Moreover, the operators $\{A_{k}\}$ have the following properties:
  1. 1.
    ${\|A_{k}\|}_{\infty}\leq e^{6(k+\ell)}\cdot{\|A\|}_{\infty}$.
  2. 2.
    $\operatorname{supp}(A_{k})\subseteq\mathrm{Ball}_{k}(\operatorname{supp}(A))$, where $\mathrm{Ball}_{k}(X)$ denotes a ball of radius $k$ around a subset of vertices $X$ under the graph’s metric.
  3. 3.
    $A_{k}$ can be calculated classically in time $T=e^{O(|\operatorname{supp}(A_{k})|)}$.
An easy corollary from the above theorem is the existence of a quasi-polynomial algorithm for the approximation of ${\langle{A}\rangle}$ for sufficiently small $\epsilon$:
Corollary 3.4
If in addition to the assumption of Theorem 3.3, the graph $G$ describes a $D$-dimensional lattice, then there exists an algorithm that for $\epsilon<\epsilon_{0}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}1/e^{6}$, an observable $A$ and an error parameter $\delta>0$, calculates ${\langle{A}\rangle}$ up to an additive error of $\delta\cdot e^{6\ell}\cdot{\|A\|}_{\infty}$, and runs in time $T=e^{O(\ell\cdot\log^{D}(1/\delta))}$. In particular, for 1D systems, the algorithm is polynomial in $1/\delta$.
Finally, we combine the results of Theorem 3.3 with the techniques used in the proof of Theorem 3.1 to show that for sufficiently small $\epsilon$, steady-state correlations between local regions (such as $2$-point correlation functions) decay exponentially with the graph distance.
Theorem 3.5
Under the same conditions of Theorem 3.3, let $\epsilon<\epsilon_{0}=1/e^{6}$ and let $\rho_{\infty}(\epsilon)$ be the steady-state of the channel $\mathcal{E}_{\epsilon}$. Then for any operators $A$ and $B$ supported in sub-regions $\operatorname{supp}(A)$ and $\operatorname{supp}(B)$ respectively, the steady-state covariance function $C(A,B)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}{\langle{AB}\rangle}-{\langle{A}\rangle}\cdot{\langle{B}\rangle}$ satisfies
$|C(A,B)|\leq c\cdot e^{6(|\operatorname{supp}(A)|+|\operatorname{supp}(B)|)}{\|A\|}\cdot{\|B\|}\cdot(\epsilon/\epsilon_{0})^{d_{AB}},$ (8)
where $c=\frac{d_{AB}+2}{(1-\epsilon/\epsilon_{0})^{2}}$ and $d_{AB}$ is the lattice distance between $\operatorname{supp}(A)$ and $\operatorname{supp}(B)$.
A detailed discussion on the results and assumptions is given in Sec. 6.
4 Proofs of the main theorems
4.1 Proof of Theorem 3.1
To prove Theorem 3.1, we draw inspiration from the multivariate expansion used to bound the AGSP Schmidt rank in Ref. [38]. In both cases, the analysis becomes much easier when moving to a multivariate setup: we replace the parameter $\epsilon$ by multiple parameters $\{\epsilon_{e}\}_{e}$, one for each edge $e\in E$. Note that we recover the original channel by setting $\epsilon_{1}=\epsilon_{2}=\ldots=\epsilon$. We use a compact notation ${\underline{\epsilon}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}(\epsilon_{1},\ldots,\epsilon_{|E|})$ to describe these parameters, $|E|$ being the total number of edges. Then our channel becomes:
$\mathcal{E}_{{\underline{\epsilon}}}(\rho)=\sum_{e=(i,j)\in E}p_{e}\Big{(}\frac{1}{2}(1-\epsilon_{e})[\mathcal{D}_{i}(\rho)+\mathcal{D}_{j}(\rho)]+\epsilon_{e}\mathcal{F}_{e}(\rho)\Big{)}.$ (9)
To proceed, we would like to expand the steady state $\rho_{\infty}({\underline{\epsilon}})$ and its corresponding Gibbs Hamiltonian in terms of parameters in ${\underline{\epsilon}}$. To that aim, we let ${\underline{e}}_{k}$ denote a subset of $k$ edges, ${\underline{e}}_{k}=\{e_{1},e_{2},\ldots,e_{k}\}$, possibly with repetition. Note that the order in which the edges appear in ${\underline{e}}_{k}$ does not matter. Given a ${\underline{e}}_{k}$, we define a shorthand notation to the product of all the $\epsilon_{e}$ parameters that correspond to the edges in ${\underline{e}}_{k}$: ${\underline{\epsilon}}^{{\underline{e}}_{k}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\epsilon_{e_{1}}\cdot\ldots\cdot\epsilon_{e_{k}}$. With this notation, a multivariable expansion of the steady state and its Gibbs Hamiltonian can be formally written as:
$\rho_{\infty}({\underline{\epsilon}})=\rho_{0}+\sum_{{\underline{e}}_{1}}{\underline{\epsilon}}^{{\underline{e}}_{1}}\rho^{({\underline{e}}_{1})}+\sum_{{\underline{e}}_{2}}{\underline{\epsilon}}^{{\underline{e}}_{2}}\rho^{({\underline{e}}_{2})}+\ldots$ (10)
$H_{G}({\underline{\epsilon}})=H_{0}+\sum_{{\underline{e}}_{1}}{\underline{\epsilon}}^{{\underline{e}}_{1}}h^{({\underline{e}}_{1})}+\sum_{{\underline{e}}_{2}}{\underline{\epsilon}}^{{\underline{e}}_{2}}h^{({\underline{e}}_{2})}+\ldots.$ (11)
Noting that $H_{G}(\epsilon)$ is obtained from $H_{G}({\underline{\epsilon}})$ by setting $\epsilon_{1}=\epsilon_{2}=\ldots=\epsilon$, we conclude that in order to prove the theorem, it is sufficient to show that
  1. 1.
    The channel $\mathcal{E}_{\underline{\epsilon}}$ has a unique fixed point $\rho_{\infty}({\underline{\epsilon}})$ for every ${\underline{\epsilon}}\in[0,1)^{n}$, and this fixed point is analytic in ${\underline{\epsilon}}$ in an open set containing $[0,1)^{n}$. Consequently, the formal expansions in (10, 11) converge.
  2. 2.
    Given a set ${\underline{e}}_{k}=\{e_{1},e_{2},\ldots,e_{k}\}$, the corresponding operator $h^{({\underline{e}}_{k})}$ is supported on the qubits that belong to the edges in ${\underline{e}}_{k}$.
  3. 3.
    $h^{({\underline{e}}_{k})}\neq 0$ only when its support forms a connected set of vertices on $G$ (and therefore it is supported on at most $k+1$ qubits).
The proof of point 1 is given by the following lemma, which is proved in Appendix A.
Lemma 4.1
Let $\mathcal{E}_{{\underline{\epsilon}}}$ be the channel from Theorem 3.1. Then the following holds:
  1. 1.
    For any ${\underline{\epsilon}}\in[0,1)^{n}$, the channel $\mathcal{E}_{{\underline{\epsilon}}}$ has a unique, full-rank steady-state $\rho_{\infty}({\underline{\epsilon}})$, and every other eigenvalue $\lambda\in\operatorname*{Spec}(\mathcal{E}_{\underline{\epsilon}})$ satisfies $|\lambda|<1$.
  2. 2.
    The matrix entries of the steady state $\rho_{\infty}({\underline{\epsilon}})$ in the computational basis are analytic functions of ${\underline{\epsilon}}$ in an open set containing $[0,1)^{n}$.
To prove points 2,3, we fix a set ${\underline{e}}_{k}=\{e_{1},e_{2},\ldots,e_{k}\}$, and consider an auxiliary graph $\tilde{G}=(V,\tilde{E})$ that is obtained from $G$ by removing all the edges that are not in ${\underline{e}}_{k}$. Let $L$ be the number of connected components in $\tilde{G}$ and let $G_{\ell}=(V_{\ell},E_{\ell})$, $\ell=1,\ldots,L$ denote the sub-graphs of $\tilde{G}$ corresponding to these connected components. Note that in some subgraphs, $E_{\ell}$ might be the empty set (when $V_{\ell}$ contains a single vertex). Next, denote by $\tilde{{\underline{\epsilon}}}$ the set of parameters obtained from ${\underline{\epsilon}}$ after setting $\tilde{\epsilon}_{e}=0$ for all $e\notin{\underline{e}}_{k}$, and consider the channel $\mathcal{E}_{\tilde{\underline{\epsilon}}}$. Such channel can be written as a sum of channels with non-overlapping supports, which correspond to the connected components of $\tilde{G}$:
$\mathcal{E}_{\tilde{\underline{\epsilon}}}=\sum_{\ell=1}^{L}P_{\ell}\mathcal{E}_{\ell}.$ (12)
Above, $\mathcal{E}_{\ell}$ is a channel acting on the qubits in $V_{\ell}$ and $\{P_{\ell}\}_{\ell}$ is a probability distribution. In such case, the steady state $\rho_{\infty}(\tilde{\underline{\epsilon}})$ decomposes into a product state
$\rho_{\infty}(\tilde{\underline{\epsilon}})=\bigotimes_{\ell=1}^{L}\sigma_{\ell}(\tilde{\underline{\epsilon}}_{\ell}),$ (13)
where $\tilde{\underline{\epsilon}}_{\ell}$ is the set of parameters in $G_{\ell}$, and each $\sigma_{\ell}(\tilde{{\underline{\epsilon}}}_{\ell})$ is the steady state of $\mathcal{E}_{\ell}$. Therefore, the corresponding Gibbs Hamiltonian is
$H_{\tilde{G}}(\tilde{{\underline{\epsilon}}})=\sum_{\ell}H_{\ell}(\tilde{{\underline{\epsilon}}}_{\ell}),\qquad H_{\ell}(\tilde{{\underline{\epsilon}}}_{\ell})=-\log\sigma_{\ell}(\tilde{{\underline{\epsilon}}}_{\ell}).$ (14)
We now make the simple observation that the perturbation expansion coefficients $\{h^{({\underline{e}}_{j})}\}$ of the Gibbs Hamiltonian $H_{G}({\underline{\epsilon}})$ in Eq. (11) do not depend on the perturbation parameters $\epsilon_{1},\epsilon_{2},\ldots$. Therefore, the same coefficients must appear in the perturbation expansion of $H_{\tilde{G}}(\tilde{{\underline{\epsilon}}})$ in Eq. (14), as it is obtained from $H_{G}({\underline{\epsilon}})$ by setting $\epsilon_{e}=0$ for all $\epsilon\notin{\underline{e}}_{k}$. In particular, as ${\underline{\epsilon}}^{{\underline{e}}_{k}}=\tilde{{\underline{\epsilon}}}^{{\underline{e}}_{k}}$, then $H_{G}(\tilde{{\underline{\epsilon}}})$ must contain the term ${\underline{\epsilon}}^{{\underline{e}}_{k}}h^{({\underline{e}}_{k})}$. Assume first that ${\underline{e}}_{k}$ does not correspond to a single connected component, i.e its edges are scattered at different connected components. There will be no single monomial $\tilde{{\underline{\epsilon}}}_{\ell}$ in Eq. (14) that contains all the variables in ${\underline{\epsilon}}^{{\underline{e}}_{k}}$, and therefore $H_{\tilde{G}}(\tilde{{\underline{\epsilon}}})$ from Eq. (14) will not contain a term that is proportional to ${\underline{\epsilon}}^{{\underline{e}}_{k}}$. This necessarily implies that $h^{({\underline{e}}_{k})}=0$, proving property (3). On the other hand, if ${\underline{e}}_{k}$ forms a connected component, then there is a subgraph $G_{\ell}$ where ${\underline{e}}_{k}\subseteq E_{\ell}$, and since by construction, $\tilde{G}$ contains only the edges in ${\underline{e}}_{k}$, we conclude that ${\underline{e}}_{k}=E_{\ell}$. Consequently, the support of $h^{({\underline{e}}_{k})}$ is contained in the support of $H_{\ell}=-\log\sigma_{\ell}$, which is the support of the qubits in ${\underline{e}}_{k}$. This proves property (2).
4.2 Proof of Theorem 3.3
The proof of Theorem 3.3 uses a rather direct perturbation theory approach. However, to obtain tighter bounds, we work in a product basis of operators that diagonalizes some of the perturbation equations. In addition, we use this basis to define an operator norm that simplifies many of the bounds. We therefore begin by defining the operator basis and norm and discussing few of their basic properties, before we give the full proof.
4.2.1 The Wauli basis
We first note that without loss of generality, we may assume that the $W_{i}$ states are diagonal in the computational basis, since we can always locally redefine the computational basis of each qubit so that it diagonalizes $W_{i}$. Moreover, using the same trick, we further assume that the eigenvector of $W_{i}$ with the largest eigenvalue is ${|{0}\rangle}$. We therefore define a set of $\lambda_{i}\in[0,1)$ such
$W_{i}=\frac{1}{2}(\mathbbm{1}+\lambda_{i}Z).$ (15)
We then define $\lambda$ to be the maximum of all these weights:
$\lambda\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max_{i}\lambda_{i}<1.$ (16)
Next, we use the $W_{i}$ states to define an operator basis that diagonalizes the $\mathcal{D}_{i}$ dissipator channel. We call it the “Wauli basis”, and it is defined as follows:
Definition 4.2 (The Wauli basis)
For a given qubit $i$, the Wauli basis is the set of operators
$\{Q^{(i)}_{0},Q^{(i)}_{1},Q^{(i)}_{2},Q^{(i)}_{3}\}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\{W_{i},X/2,Y/2,Z/2\}.$ (17)
In order to keep the notation clean, we shall drop the $(i)$ index from the basis, where it is clear context on which qubit we act. The Wauli basis can be naturally extended to a multiqubit setup. For a string ${\underline{\smash{\alpha}}}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{n})$ with $\alpha_{i}\in\{0,1,2,3\}$, define $Q_{\underline{\smash{\alpha}}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}Q_{\alpha_{1}}\otimes Q_{\alpha_{2}}\otimes\ldots\otimes Q_{\alpha_{n}}$. We call $\{Q_{\underline{\smash{\alpha}}}\}_{{\underline{\smash{\alpha}}}}$ the Wauli basis of the $n$-qubits space. With a slight abuse of notation, we shall say that the support of the string ${\underline{\smash{\alpha}}}$ is the set of sites in which ${\underline{\smash{\alpha}}}_{i}\neq 0$, and write $|{\underline{\smash{\alpha}}}|$ to represent the size of the support.
Note that the Wauli basis elements are normalized with respect to the trace norm, i.e., ${\|Q_{\alpha}\|}_{1}=1$. Also note that they diagonalize the dissipators: indeed, under the settings of Theorem 3.3, $\mathcal{D}_{i}$ is given by $\mathcal{D}_{i}(\rho)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathrm{Tr}_{i}(\rho)\otimes W_{i}$, and therefore $\mathcal{D}_{i}(Q^{(i)}_{0})=Q^{(i)}_{0}$ and $\mathcal{D}_{i}(Q^{(i)}_{\alpha})=0$ for $\alpha=1,2,3$. Finally, we see that while the Wauli basis is not orthonormal with respect to the Hilbert-Schmidt inner product, it is still a valid basis, as its elements are linearly independent. Moreover, we can still make use of orthogonality by looking at its dual basis $\{\tilde{Q}_{\underline{\smash{\beta}}}\}_{{\underline{\smash{\beta}}}}$ that is defined by the equations
${\langle{\tilde{Q}_{\underline{\smash{\beta}}},Q_{\underline{\smash{\alpha}}}}\rangle}_{HS}=\delta_{{\underline{\smash{\alpha}}},{\underline{\smash{\beta}}}},\quad\forall{\underline{\smash{\alpha}}},{\underline{\smash{\beta}}}.$ (18)
Explicitly, this basis is given by
Definition 4.3 (The Dual-Wauli basis)
For the qubit $i$, the Dual Wauli basis is given by
$\{\tilde{Q}^{(i)}_{0},\tilde{Q}^{(i)}_{1},\tilde{Q}^{(i)}_{2},\tilde{Q}^{(i)}_{3}\}=\{\mathbbm{1},X,Y,Z-\lambda_{i}\mathbbm{1})\}.$ (19)
As in the Wauli basis case, we will often drop the $(i)$ index and expand this basis to multiple qubits by setting $\tilde{Q}_{\underline{\smash{\alpha}}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\tilde{Q}_{\alpha_{1}}\otimes\tilde{Q}_{\alpha_{2}}\otimes\ldots\otimes\tilde{Q}_{\alpha_{n}}$. Note that as $\tilde{Q}^{(i)}_{0}=\mathbbm{1}$, then for any string ${\underline{\smash{\alpha}}}$, it holds that $|{\underline{\smash{\alpha}}}|=|\operatorname{supp}(\tilde{Q}_{\underline{\smash{\alpha}}})|$.
Just as the Wauli basis diagonalizes the Dissipator $\mathcal{D}_{i}$ channel, its dual diagonalizes its adjoint map $\mathcal{D}^{*}_{i}$ that is given by $\mathcal{D}^{*}_{i}(A)=\mathrm{Tr}_{i}(AW_{i})\otimes\mathbbm{1}_{i}$.
Throughout our proof we shall work in the Heisenberg picture, where we apply a series of adjoint channels on $A$. Consequently, it will be beneficial to work with the dual Wauli basis, which diagonalizes the adjoint dissipators $\mathcal{D}_{i}^{*}$. In addition, we shall use the following operator norm, which is defined by the dual Wauli basis:
Definition 4.4 ($Q_{1}$ norm)
Given an operator $O=\sum_{\underline{\smash{\alpha}}}c_{\underline{\smash{\alpha}}}\tilde{Q}_{\underline{\smash{\alpha}}}$ expanded in the dual-Wauli basis, we define its $Q_{1}$ norm to be the $L_{1}$ norm of its coefficients vector, i.e.,
${\|O\|}_{Q_{1}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{{\underline{\smash{\alpha}}}}|c_{\underline{\smash{\alpha}}}|.$ (20)
We shall also need the following definition subspaces of operators with a bounded support with respect to the dual Wauli basis:
Definition 4.5 (The $\mathcal{S}_{k}$ subspaces)
For any integer $k\geq 0$, we let $\mathcal{S}_{k}$ denote the set of operators spanned by $\tilde{Q}_{\underline{\smash{\alpha}}}$ with $|{\underline{\smash{\alpha}}}|\leq k$ (i.e., the set of operators spanned by dual Wauli basis with support $\leq k$).
Using the definitions above, the $Q_{1}$ norm satisfies three additional properties that are described in the following lemma:
Lemma 4.6
  1. 1.
    For any operator $O$ on $n$ qubits and the identity operator $\mathbbm{1}$ on $n^{\prime}$ qubits,
    ${\|O\otimes\mathbbm{1}\|}_{Q_{1}}={\|O\|}_{Q_{1}}$ (21)
  2. 2.
    For any $O\in\mathcal{S}_{k}$,
    ${\|O\|}_{\infty}$ $\leq(1+\lambda)^{k}\cdot{\|O\|}_{Q_{1}}.$ (22)
  3. 3.
    For every $O$ supported on $\ell$ qubits,
    ${\|O\|}_{Q_{1}}$ $\leq 4^{\ell}\cdot{\|O\|}_{\infty}.$ (23)
Finally, we shall need the following lemma, which upperbounds the growth in the $Q_{1}$ norm due to the action of the adjoint of a $k$-local channel:
Lemma 4.7
Let $\mathcal{E}$ be a channel acting on $k$ qubits, and let $\mathcal{E}^{*}$ be its adjoint. Then for any operator $O$,
${\|\mathcal{E}^{*}(O)\|}_{Q_{1}}\leq 4^{k}(1+\lambda)^{k}\cdot{\|O\|}_{Q_{1}}\leq 2^{3k}\cdot{\|O\|}_{Q_{1}}.$ (24)
The proofs of both lemmas can be found in Appendix B. We are now finally in position to prove Theorem 3.3.
4.2.2 Proof of Theorem 3.3
We start by rewriting the channel in Eq. (3) as
$\mathcal{E}_{\epsilon}(\rho)=(1-\epsilon)\sum_{i\in V}q_{i}\mathcal{D}_{i}+\epsilon\sum_{e\in E}p_{e}\mathcal{F}_{e}(\rho),$ (25)
where $q_{i}=\frac{1}{2}\sum_{e\ni i}p_{e}$ is a probability distribution on the vertices. We then decompose
$\mathcal{E}_{\epsilon}=\mathcal{E}_{0}+\epsilon\mathcal{E}_{1},$
where
$\mathcal{E}_{0}(\rho)$ =$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i\in V}q_{i}\mathcal{D}_{i}(\rho)=\sum_{i\in V}q_{i}\mathrm{Tr}_{i}(\rho)\otimes W_{i},$ (26)
$\mathcal{E}_{1}(\rho)$ =$\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{e\in E}p_{e}\mathcal{F}_{e}(\rho)-\sum_{i\in V}q_{i}\mathrm{Tr}_{i}(\rho)\otimes W_{i}.$ (27)
Plugging the expansion $\rho_{\infty}(\epsilon)=\rho_{0}+\epsilon\rho_{1}+\epsilon^{2}\rho_{2}+\ldots$ into the steady-state equation $\mathcal{E}_{\epsilon}(\rho_{\infty})=\rho_{\infty}$, we obtain the steady-state perturbation equations
$\mathcal{E}_{0}(\rho_{0})$ $=\rho_{0},$ (28)
$\rho_{k}-\mathcal{E}_{0}(\rho_{k})$ $=\mathcal{E}_{1}(\rho_{k-1}),\quad k=1,2,3\ldots$ (29)
The first equation can be solved directly, yielding:
$\rho_{0}=\bigotimes_{i}\rho_{0}^{(i)}=\bigotimes_{i}W_{i}.$ (30)
For higher order $\rho_{k}$, note that both sides of Eq. (29) are defined by trace killing maps, and therefore they do not determine the part in $\rho_{k}$ that is proportional to $\mathbbm{1}$. However, since $\mathrm{Tr}(\rho_{0})=1$ and the full $\rho_{\infty}(\epsilon)$ must have unit trace, then clearly $\mathrm{Tr}(\rho_{k})=0$ for all $k>0$, which fixes the $\mathbbm{1}$ contribution in each $\rho_{k}$. Therefore, we may define a new super-operator $\mathcal{K}$
$\mathcal{K}(\rho_{k})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\rho_{k}-\mathcal{E}_{0}(\rho_{k})+\mathrm{Tr}(\rho_{k})\rho_{0}$ (31)
and replace Eq. (29) with
$\mathcal{K}(\rho_{k})=\mathcal{E}_{1}(\rho_{k-1}).$
The advantage of working with this equation is that, as we will show shortly, $\mathcal{K}$ is an invertible map. Therefore, it gives us recursive expression for $\rho_{k}$:
$\rho_{k}=(\mathcal{K}^{-1}\circ\mathcal{E}_{1})(\rho_{k-1}).$ (32)
Next, consider our local observable $A$, whose expectation value at the steady state can be written as
${\langle{A}\rangle}=\mathrm{Tr}\big{(}A\rho_{\infty}(\epsilon))=\mathrm{Tr}(A\rho_{0})+\epsilon\mathrm{Tr}(A\rho_{1})+\ldots$
We can use the recursive identity Eq. (32) and move the super-operators to the Heisenberg picture, obtaining:
$\mathrm{Tr}(A\rho_{k})=\mathrm{Tr}\big{(}A\cdot(\mathcal{K}^{-1}\circ\mathcal{E}_{1})(\rho_{k-1})\big{)}=\mathrm{Tr}\big{(}(\mathcal{K}^{-1}\circ\mathcal{E}_{1})^{*}(A)\cdot\rho_{k-1}\big{)},\qquad k=1,2,\ldots$ (33)
Therefore, defining the transition super-operator $\mathcal{T}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\mathcal{E}_{1}^{*}\circ(\mathcal{K}^{*})^{-1}$, we get the recursive equation
$\mathrm{Tr}(A\rho_{k})=\mathrm{Tr}\big{(}\mathcal{T}(A)\rho_{k-1}\big{)},$
which lets us define a series of operators $\{A=A_{0},A_{1},A_{2},\ldots,A_{k}\}$ such that
$A_{k}=\mathcal{T}^{k}(A),\qquad\mathrm{Tr}(A_{k}\rho_{0})=\mathrm{Tr}(A\rho_{k}).$ (34)
This proves Eq. (7).
For the next part of the proof, we will find an upper-bound to ${\|A_{k}\|}_{\infty}$ by working in the dual Wauli basis, and see how it transforms under $\mathcal{T}$. Firstly, from (26) and (27), we get
$\mathcal{E}^{*}_{0}(A)$ $=\sum_{i\in V}q_{i}\mathrm{Tr}(AW_{i})\otimes\mathbbm{1}_{i},$ (35)
$\mathcal{E}_{1}^{*}(A)$ $=\sum_{e\in E}p_{e}\mathcal{F}_{e}^{*}(A)-\mathcal{E}_{0}^{*}(A).$ (36)
Notice that $\{Q_{\underline{\smash{\alpha}}}\}_{\underline{\smash{\alpha}}}$ and $\{\tilde{Q}_{\underline{\smash{\alpha}}}\}_{\underline{\smash{\alpha}}}$ are the eigenbases of $\mathcal{E}_{0}$ and $\mathcal{E}_{0}^{*}$ respectively, with eigenvalues:
$\mathcal{E}_{0}(Q_{\underline{\smash{\alpha}}})=\sum_{i\notin\operatorname{supp}({\underline{\smash{\alpha}}})}q_{i}Q_{\underline{\smash{\alpha}}}=(1-q_{{\underline{\smash{\alpha}}}})Q_{\underline{\smash{\alpha}}},\qquad\mathcal{E}_{0}^{*}(\tilde{Q}_{\underline{\smash{\alpha}}})=(1-q_{{\underline{\smash{\alpha}}}})\tilde{Q}_{\underline{\smash{\alpha}}},$ (37)
where we defined $q_{{\underline{\smash{\alpha}}}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i\in\operatorname{supp}({\underline{\smash{\alpha}}})}q_{i}$ and used the fact that $\{q_{\underline{\smash{\alpha}}}\}$ form a probability distribution. Using the above equations, together with the definition of $\mathcal{K}$ in Eq. (31) and the fact that $\mathrm{Tr}(Q_{\underline{\smash{\alpha}}})=\delta_{{\underline{\smash{\alpha}}},{\underline{0}}}$ where 0¯=(${\underline{0}}=(0,\ldots,0)$, and that $\rho_{0}=Q_{\underline{0}}$ (see Eq. (30)), we conclude that
$\mathcal{K}(Q_{\underline{\smash{\alpha}}})=\begin{cases}q_{\underline{\smash{\alpha}}}Q_{\underline{\smash{\alpha}}},&{\underline{\smash{\alpha}}}\neq{\underline{0}}\\ Q_{\underline{\smash{\alpha}}},&{\underline{\smash{\alpha}}}={\underline{0}}\end{cases}.$ (38)
Consequently, $\mathcal{K}$ is diagonal in the Wauli basis, and as its eigenvalues are non-zero, it is also invertible. A parallel argument applies for its conjugate $\mathcal{K}^{*}$ with respect to the dual Wauli basis, and therefore, also $\mathcal{K}^{*}$ is invertible and
$(\mathcal{K}^{*})^{-1}(\tilde{Q}_{\underline{\smash{\alpha}}})=\begin{cases}\frac{1}{q_{\underline{\smash{\alpha}}}}\tilde{Q}_{\underline{\smash{\alpha}}},&{\underline{\smash{\alpha}}}\neq{\underline{0}}\\ \tilde{Q}_{\underline{\smash{\alpha}}},&{\underline{\smash{\alpha}}}={\underline{0}}\end{cases}.$ (39)
Recalling that $\mathcal{T}=\mathcal{E}^{*}_{1}\circ(\mathcal{K}^{*})^{-1}$, we can now proceed to find how it acts on the dual Wauli basis. This is done in the following lemma, proven in Appendix B:
Claim 4.8
  1. 1.
    Recall the definition of the $\mathcal{S}_{k}$ subspace in Def. 4.5. Then for every $O\in\mathcal{S}_{k}$, we have $\mathcal{T}(O)\in\mathcal{S}_{k+1}$. Moreover, the support of $\mathcal{T}(O)$ is included in the ball of radius $1$ around the support of $O$.
  2. 2.
    For any observable $O$, it holds ${\|\mathcal{T}(O)\|}_{Q_{1}}\leq 130\cdot{\|O\|}_{Q_{1}}$.
Let us use the claim to establish the upper bound on ${\|A_{k}\|}_{\infty}$. As $A_{k}=\mathcal{T}^{k}(A)$, then by Bullet 2, we get:
${\|A_{k}\|}_{Q_{1}}\leq 130{\|A_{k-1}\|}_{Q_{1}}\leq\ldots\leq 130^{k}{\|A\|}_{Q_{1}}.$ (40)
Next, we use Lemma 4.6 to transform the above inequality into a similar inequality in the operator norm. By assumption, $A$ is supported on $\ell$ qubits, and therefore by Bullet 1 of Claim 4.8, $A_{k}=\mathcal{T}^{k}(A)\in\mathcal{S}_{k+\ell}$. Using inequality (22) from Lemma 4.6, we get
${\|A_{k}\|}_{\infty}\leq(1+\lambda)^{k+\ell}\cdot{\|A_{k}\|}_{Q_{1}}.$ (41)
On the other hand, by inequality (23) of the same lemma, combined with the fact that $|\operatorname{supp}(A)|=\ell$, we get
${\|A\|}_{Q_{1}}\leq 4^{\ell}\cdot{\|A\|}_{\infty}.$ (42)
Altogether, inequalities (4142) imply that
${\|A_{k}\|}_{\infty}\leq\big{(}130(1+\lambda)\big{)}^{k}\cdot\big{(}4(1+\lambda)\big{)}^{\ell}\cdot{\|A\|}_{\infty}\leq e^{6(k+\ell)}\cdot{\|A\|}_{\infty}.$ (43)
To conclude the proof, let us estimate $|\operatorname{supp}(A_{k})|$ and the complexity of calculating $A_{k}$. To do that, consider the how $\mathcal{T}(\tilde{Q}_{\underline{\smash{\alpha}}})$ changes the support of $\tilde{Q}_{\underline{\smash{\alpha}}}$. By Eq. (39) and the definition of $\mathcal{T}$, we know that
$\mathcal{T}(\tilde{Q}_{\underline{\smash{\alpha}}})=\mathcal{E}_{1}^{*}\circ(\mathcal{K}^{*})^{-1}(\tilde{Q}_{\underline{\smash{\alpha}}})\propto\mathcal{E}_{1}^{*}(\tilde{Q}_{\underline{\smash{\alpha}}}).$
From the expression of $\mathcal{E}_{1}^{*}$ in Eq. (36), we see that as $\mathcal{F}_{e}^{*}(\mathbbm{1})=\mathcal{E}_{0}^{*}(\mathbbm{1})=\mathbbm{1}$ (which is the completeness property of quantum channels), then the only correlators $\mathcal{F}_{e}^{*}$ acting non-trivially on $\tilde{Q}_{\underline{\smash{\alpha}}}$ correspond to edges that intersect with $\operatorname{supp}({\underline{\smash{\alpha}}})$. Therefore, we can write
$\mathcal{E}_{1}^{*}(\tilde{Q}_{\underline{\smash{\alpha}}})=\sum_{\underline{\smash{\beta}}}c_{\underline{\smash{\beta}}}\tilde{Q}_{\underline{\smash{\beta}}},$
where $c_{\underline{\smash{\beta}}}\neq 0$ only when $\operatorname{supp}(\tilde{Q}_{\underline{\smash{\beta}}})$ can be obtained from $\operatorname{supp}(\tilde{Q}_{\underline{\smash{\alpha}}})$ by adding a neighboring qubit (with respect to $G$) to one of the qubits in $\operatorname{supp}(\tilde{Q}_{\underline{\smash{\alpha}}})$. Geometrically, this can be viewed as if under $\mathcal{E}_{1}^{*}$ (and therefore also under $\mathcal{T}$), the support of $\tilde{Q}_{\underline{\smash{\alpha}}}$ spreads by one step over the graph to every direction. This implies that $\operatorname{supp}(A_{k})$ is included in the union of all $k$-walks over $G$ that started from the support of $A=A_{0}$. When the underlying graph $G$ represents a $D$-dimensional lattice, then $\operatorname{supp}(A_{k})$ grows like $O(\ell\cdot k^{D})$.
To estimate the complexity of calculating $A_{k}$, we perform the calculation in the dual Wauli basis. The number of basis elements needed to represent $A_{k}$ is $\mathcal{O}(4^{|\operatorname{supp}(A_{k})|})$, and therefore since $A_{k}=\mathcal{T}(A_{k-1})$, the calculation of $A_{k}$ can be reduced to the a matrix-vector multiplication over a linear space of dimension $q=4^{|\operatorname{supp}(A_{k})|}$, which can be done in time $O(q^{2})=O(4^{2|\operatorname{supp}(A_{k})|})=e^{O(|\operatorname{supp}(A_{k})|)}$. This is also an upper bound for the first $k-1$ steps of the calculation, and therefore the total running time needed to calculate $A_{k}$ is $e^{O(|\operatorname{supp}(A_{k})|)}$.
4.3 Proof of Theorem 3.5
The idea behind the proof is as follows. We expand the covariance function as a power series in $\epsilon$, and show that the lowest non-vanishing contribution comes from terms of degree larger than the graph distance between the supports of $A$ and $B$. Theorem 3.3 will guarantee that these high order terms decay exponentially. To show the vanishing of the low-order terms, we move the multivariate settings used in the proof of Theorem 3.1, where we assign different $\epsilon_{j}$ to different edge interactions. Expanding the covariance in powers of $\{\epsilon_{j}\}_{j}$, it follows that non-vanishing terms arise from sequences $\epsilon$’s that connect $A$ to $B$. This shows that non-zero contribution will come from monomials of degree which is larger than the distance between $A$ and $B$. These, however, drop exponentially in $\epsilon/\epsilon_{0}$ due to bullet 1 of Theorem 3.3.
Let us begin by considering the channel from Theorem 3.3, namely $\mathcal{E}_{\epsilon}=\sum_{e\in E}p_{e}\mathcal{E}_{\epsilon}^{(e)}$, and assign different variable $\epsilon_{e}$ to each edge $\mathcal{E}_{e}$. The modified channel becomes $\sum_{e\in E}p_{e}\mathcal{E}_{\epsilon_{e}}^{(e)}$ as presented in Eq. (9), whose unique steady state $\rho_{\infty}({\underline{\epsilon}})$ (given in Eq. (10)) is analytic in ${\underline{\epsilon}}$, as claimed in Lemma 4.1. As a result, the steady state covariance is analytic, which allows us to expand it in the following manner:
$C(A,B)_{\rho_{\infty}({\underline{\epsilon}})}=C_{0}+\sum_{{\underline{e}}_{1}}{\underline{\epsilon}}^{{\underline{e}}_{1}}C^{({\underline{e}}_{1})}+\sum_{{\underline{e}}_{2}}{\underline{\epsilon}}^{{\underline{e}}_{2}}C^{({\underline{e}}_{2})}+\ldots$ (44)
Above, we used the same notation ${\underline{e}}_{k}=\{e_{1},\ldots,e_{k}\}$, ${\underline{\epsilon}}^{{\underline{e}}_{k}}=\epsilon_{e_{1}}\dots\epsilon_{e_{k}}$ that we used in Eq. (10), and $\{C^{({\underline{e}}_{k})}\}$ are scalar coefficients. Note that $C_{0}=0$, because when setting $\epsilon=0$, the steady state becomes a product state and the covariance vanishes.
We start by showing that the expansion consists only of terms that connect $A$ to $B$. Fix ${\underline{e}}_{k}=\{e_{1},\ldots,e_{k}\}$, and assume it does not contain a path from $A$ to $B$. As in Sec. 4.1, define $\tilde{G}$ to be the ${\underline{e}}_{k}$-induced subgraph, which is achieved from $G$ by restricting to the edges in ${\underline{e}}_{k}$. Note that due to the assumption on ${\underline{e}}_{k}$, each of the disjoint connected components of $\tilde{G}$ intersect with $A$ or $B$ or none of them, but not $A$ and $B$ simultaneously. Next, as done in the proof of Theorem 3.1, we denote by $\tilde{{\underline{\epsilon}}}$ the set of parameters obtained from ${\underline{\epsilon}}$ after setting $\tilde{\epsilon}_{e}=0$ for all $e\notin{\underline{e}}_{k}$, and consider the reduced channel $\mathcal{E}_{\tilde{\underline{\epsilon}}}=\sum_{\ell=1}^{L}P_{\ell}\mathcal{E}_{\ell}$, which is a convex sum of disjoint channels on the components of $\tilde{G}$ (see Eq. (12)). Then corresponding steady state $\rho_{\infty}(\tilde{\underline{\epsilon}})$ decomposes to a tensor product of states,
$\rho_{\infty}(\tilde{\underline{\epsilon}})=\bigotimes_{\ell=1}^{L}\sigma_{\ell}(\tilde{\underline{\epsilon}}_{\ell})=\sigma_{G_{A}}\otimes\sigma_{G_{B}}\otimes\sigma_{other},$ (45)
where $G_{A}$ ($G_{B}$) is the union of components of $\tilde{G}$ that intersect with $A$ ($B$). This is because ${\underline{e}}_{k}$ does not contain a path from $A$ to $B$, $G_{A}$ and $G_{B}$ are disjoint, which guarantees that $\rho_{\infty}(\tilde{\underline{\epsilon}})$ is indeed a product state. Moreover, the $\rho_{\infty}(\tilde{\underline{\epsilon}})$-covariance function satisfies
$C(O_{A},O_{B})_{\rho_{\infty}(\tilde{\underline{\epsilon}})}=\mathrm{Tr}[O_{A}O_{B}\cdot\sigma_{G_{A}}\otimes\sigma_{G_{B}}]-\mathrm{Tr}[O_{A}\sigma_{G_{A}}]\cdot\mathrm{Tr}[O_{B}\sigma_{G_{B}}]=0.$ (46)
Expanding the expression above in $\tilde{\underline{\epsilon}}$, its vanishing, together with analyticity implies that each monomial vanishes individually, and in particular, $C^{({\underline{e}}_{k})}=0$.
We conclude that the expansion $C({\underline{\epsilon}})=\sum_{\underline{e}}\epsilon^{{\underline{e}}}C^{({\underline{e}})}$ contains only terms that connect $A$ and $B$. Such terms must be of degree which is greater than or equal to the distance $d_{AB}$ between the supports of $A$ and $B$.
We now return to the original settings of the problem by setting $\epsilon_{1}=\epsilon_{2}=\dots=\epsilon$ and expand the covariance as a power series in $\epsilon$:
$C(A,B)_{\rho_{\infty}(\epsilon)}={\langle{AB}\rangle}_{\rho_{\infty}(\epsilon)}-{\langle{A}\rangle}_{\rho_{\infty}(\epsilon)}\cdot{\langle{B}\rangle}_{\rho_{\infty}(\epsilon)}=\sum_{k\geq d_{AB}}\epsilon^{k}C^{(k)}.$ (47)
Expanding ${\langle{AB}\rangle}$, ${\langle{A}\rangle}$ and ${\langle{B}\rangle}$ according to Eq. (7) of Theorem 3.3, we get the following expression for $C^{(k)}$:
$C^{(k)}=\mathrm{Tr}(\rho_{0}(AB)_{k})-\sum_{j=0}^{k}\mathrm{Tr}(\rho_{0}A_{j})\cdot\mathrm{Tr}(\rho_{0}B_{k-j}).$ (48)
Using triangle inequality and plugging Bullet 1 of Theorem 3.3 to Eq. (48), we achieve the following upper bound
$|C^{(k)}|$ $\leq|\mathrm{Tr}(\rho_{0}(AB)_{k})|+\sum_{j=0}^{k}|\mathrm{Tr}(\rho_{0}A_{j})|\cdot|\mathrm{Tr}(\rho_{0}B_{k-j})|$
$\leq e^{6(k+\ell_{A}+\ell_{B})}{\|A\otimes B\|}+\sum_{j=0}^{k}e^{6(j+\ell_{A})}{\|A\|}\cdot e^{6(k-j+\ell_{B})}{\|B\|}$
$\leq e^{6(k+\ell_{A}+\ell_{B})}{\|A\|}\cdot{\|B\|}+(k+1)e^{6(k+\ell_{A}+\ell_{B})}{\|A\|}\cdot{\|B\|}$
$=(k+2)e^{6(k+\ell_{A}+\ell_{B})}{\|A\|}\cdot{\|B\|}.$
Here, $\ell_{A}$ and $\ell_{B}$ denote the number of qubits in the support of $A$ and $B$, respectively. Plugging this to the expansion (47) produces the desired decay
$|C(A,B)_{\rho_{\infty}(\epsilon)}|\leq\sum_{k\geq d_{AB}}\epsilon^{k}|C^{(k)}|\leq{\|A\|}\cdot{\|B\|}e^{6(\ell_{A}+\ell_{B})}\sum_{k\geq d_{AB}}(k+2)\left(\frac{\epsilon}{e^{-6}}\right)^{k}.$
Computing the geometric sum $\sum_{k\geq d}(k+2)q^{k}=q^{d}\frac{(d+2)+\frac{q}{1-q}}{1-q}<q^{d}\frac{d+2}{(1-q)^{2}}$ gives Eq. (8).
5 Numerical findings
In this section, we present our numerical results of computer simulations on small systems. We considered channels in the restricted form stated in Theorem 3.3 on a 1D-ring of $n=10$ qubits:
$\mathcal{E}_{\epsilon}(\rho)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{\epsilon}{10}\sum_{i=0}^{9}\mathcal{F}_{i,i+1}(\rho)+\frac{1-\epsilon}{10}\sum_{i=0}^{9}\mathrm{Tr}_{i}(\rho)\otimes W_{i},$ (49)
where we identified $i=10$ with $i=0$ for periodic boundary conditions. The main purpose of these simulations was to examine the decay of ${\|H_{k}\|}$ (or, more precisely, of $\epsilon^{k}{\|H_{k}\|}$) as a function of $k$ for typical channels.
In all the simulations, we initialized our qubits to the state $\rho_{0}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}{|{0}\rangle\langle{0}|}^{\otimes n}$ and repeatedly applied the channel until it approximately converged to the steady state. We denoted by $\rho_{t}$ the state of the system at step $t$, and assumed convergence to the steady state $\rho_{\infty}(\epsilon)$ when ${\|\rho_{t}-\mathcal{E}(\rho_{t})\|}_{1}\leq 10^{-8}$. Once we reached the steady state, we calculated $H_{G}(\epsilon)=-\log{\rho_{\infty}(\epsilon)}$, the Gibbs Hamiltonian of our channel.
Having calculated $H_{G}(\epsilon)$ for various $\epsilon\in[0,1]$, our task was to estimate the norm of the $H_{k}$ terms in statement of Theorem 3.1. This, however, turned out to be a numerically challenging task, as one needs to perform some sort of an interpolation in $\epsilon$ in order to extract the $H_{k}$ from $H_{G}(\epsilon)$. Such procedures usually become numerically unstable as $k$ increases. Instead, we circumvented this difficulty by estimating a closely related quantity. Given $H_{G}(\epsilon)$, we expanded it in terms of Pauli strings, $H_{G}(\epsilon)=\sum_{\underline{\smash{\alpha}}}c_{\underline{\smash{\alpha}}}(\epsilon)P_{\underline{\smash{\alpha}}}$, and defined $\hat{H}_{k}(\epsilon)$ by
$\hat{H}_{k}(\epsilon)\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\frac{\sum_{d({\underline{\smash{\alpha}}})=k}c_{\underline{\smash{\alpha}}}(\epsilon)P_{\underline{\smash{\alpha}}}}{{\|\sum_{d({\underline{\smash{\alpha}}})=1}c_{\underline{\smash{\alpha}}}(\epsilon)P_{\underline{\smash{\alpha}}}\|}_{\infty}}.$ (50)
In the above formula, $d({\underline{\smash{\alpha}}})$ denotes the diameter of the support of ${\underline{\smash{\alpha}}}$. As we worked on a ring, it was calculated as follows. Given a ${\underline{\smash{\alpha}}}=(\alpha_{0},\ldots,\alpha_{n-1})$, we listed all trivial segments in the string, i.e., contiguous locations where $\alpha_{i}=0$, and defined $d({\underline{\smash{\alpha}}})$ to be $n=10$ minus the length of the longest segment. The denominator in the above definition is an overall factor that guarantees that ${\|\hat{H}_{k=1}(\epsilon)\|}_{\infty}=1$.
[…]
The first dynamics we simulated was composed of random $W_{i}$ and random correlators $\{\mathcal{F}_{e=(i,i+1)}\}$. The former where chosen randomly on the Bloch sphere, while the latter were chosen as random Kraus operators of $\mathrm{rank}=3$ (see Ref. [39] for more details). To validate the robustness of our results, we ran several random instances of such channels, and observed similar results in all runs. We plot ${\|\hat{H}_{k}(\epsilon)\|}_{\infty}$ as function of $k$ for a single instance of such channel in Fig. 1(a).
(a) First model: random Kraus maps
(b) Second model: random unitaries ($k=1)$
(c) Third model: channel from Eq. (4)
(d) Second model: random unitary ($k=2)$
Figure 1: Operator norms of $\hat{H}_{k}$ (see Eq. (50)) as function of $k$ are displayed for simulations of four choices of $\mathcal{E}_{\epsilon}$ at $\epsilon=0.1,0.2,\ldots,1$ on $10$ qubits.
The evident exponential decay of the norms for each $\epsilon=0.1,0.2,\ldots,1$ suggests that $H_{G}(\epsilon)$ is indeed quasi-local. The decay rates decrease (moderately) as $\epsilon$ is increased. We note that there is a very little difference between the decay at $\epsilon=1$ and the decay rates at $\epsilon<1$. We believe this is due to the fact that the random correlators create a channel that is still dissipative, local, and ergodic, even when the dissipators are absent.
The second channel we simulated is identical to the first, except that the correlators $\mathcal{F}_{e}$ taken to be random unitary channels: $\mathcal{F}_{e}(\rho)=\frac{1}{k}\sum_{a=1}^{k}U^{(e)}_{a}\rho U^{(e)\dagger}_{a}$, where $U^{(e)}_{a}$ are random unitaries taken from the Haar measure on two qubits, and $k=1,2$. As in the previous case, we verified robustness by running several such channels, which all yielded similar results. Results of one instance with $k=1$ are plotted in Fig. 1(b). The decays we see in Fig. 1(b) resemble the decays seen in Fig. 1(a), but there is one notable difference: Unlike in Fig. 1(a), there is a significant difference in the decay rate obtained from $\epsilon=0.9$ and from $\epsilon=1$, where we see a substantial decrease in the decay rate for $\epsilon=1$. We conjecture that in the absence of dissipators, when there is a single unitary on each edge, ergodicity is weak [^2] We suspect that for $k=1$, the number of products of unitaries required to achieve the full operator algebra is large. (see Wielandt’s inequality [37, 40]). Therefore, the channel might have eigenvalues that approach $1$, creating long range correlations within the steady state., which might lead to a slow decay in ${\|\hat{H}_{k}(\epsilon=1)\|}$. We confirm this by performing the same simulation with $k=2$, i.e., a mixture of two unitaries on each edge, which is likely to have stronger ergodicity. The results of a single instance are presented in Fig. 1(d). Note that a similar decay profile is found in $\epsilon=1$ and $\epsilon=0.9$.
The third channel we studied the specific channel that was given in Eq. (4) in Sec. 3. This is a translationally-invariant channel, in which all the $W_{i}$ are given by $W_{i}=0.9{|{+}\rangle\langle{+}|}+0.1{|{-}\rangle\langle{-}|}$ and the correlators are simply control-Z gates. The results of this simulation are plotted in Fig. 1(c). Here, at $\epsilon=1$ we see the channel leads to long-range Gibbs Hamiltonian $H_{G}$, which is not at all quasi-local. In this case the channel reduces to $\frac{1}{10}\sum_{i}\text{C-Z}_{i,i+1}\cdot\rho\cdot\text{C-Z}_{i,i+1}$ for which the steady state is not unique (in-fact, any $\rho$ that is diagonal in the computational basis is a fixed-point). For the remaining choices of $\epsilon$ values, one can see an exponential decay in agreement, which is what we expect. It would be interesting to study the long-range to short-range transition in $\epsilon$ for different choices of gates. We leave this for future research.
Overall, in our simulations we see that whenever a channel is of the form given in Theorem 3.1, we obtain a quasi-local Hamiltonian. In the limit $\epsilon=1$, the channel may or may not converge to a steady-state with quasi-local Gibbs Hamiltonian.
6 Discussion and future research
In this paper, we established rigorous statements on the steady states of a broad family of local stochastic dynamics, defined as a convex sum of local channels (3). These are channels that are composed of convex combination of single site channels (named dissipators), and two-local channels (correlators), for which the weight of the correlators is given by a parameter $\epsilon\in[0,1)$. Our first main result is given in Theorem 3.1, where we show that the steady state can be written as a Gibbs state with a Gibbs Hamiltonian $H_{G}$ that admits a series expansion in $\epsilon$. We proved that the $k$’th order in $H_{G}$ is a $(k+1)$-geometrically local Hamiltonian $H_{k}$. While we were not able to derive an explicit upper bound on the norm of each $H_{k}$, we used numerical simulations of several exemplary channels of small systems, to show that the norm of $\epsilon^{k}H_{k}$ decays exponentially in $k$ for all $\epsilon\in[0,1)$.
Although the statement of Theorem 3.1 is intuitive, as the steady state is reached by local dynamics, it is not a trivial statement, as even in the classical regime local dynamical rule might converge to a long range Gibbs Hamiltonian [41]. The proof of the theorem follows an idea also used in Ref. [38] of moving to a multivariate expansion, which gives rise to a short combinatorial proof. We note that our proof is given in the language of quantum channels, but its generalization to the more general setup of local Lindbladians is straight-forward.
Below, we give several open questions and possible research direction related to Theorem 3.1
  1. 1.
    Bounding the norm of $H_{k}$. Arguably, the biggest missing piece in Theorem 3.1 is a bound on the norm of $H_{k}$ terms. Showing that ${\|H_{k}\|}\leq\mu^{k}$ for some $\mu>0$, will rigorously imply that the Gibbs Hamiltonian is quasi-local for $\epsilon<1/\mu$. While we do not have a proof for that, we suspect that this is the case. We believe that the techniques used in the proof of Theorem 3.3, where a similar threshold was established, can be used to prove a statement like that. We leave that for future research.
  2. 2.
    Generalization to general $2$-local channels: Setting $\epsilon=1$ in (3) leads to a general 2-local stochastic dynamics, which is not expected to converge to a quasi-local Gibbs state in all cases. Nevertheless, it is often observed in the numerics that this is indeed the case. A necessary condition for this to happen is that the channel needs to be ergodic, as shown in Ref. [41], as well as in the numerical data in Fig. 1(c). One possible explanation for the convergence of general $2$-local channels is that in these cases, the general $2$-local channels can be self-decomposed into a sum of 1-local and 2-local channels. This way, an effective small parameter $\epsilon$ can be extracted from the dynamics itself.
  3. 3.
    Engineering Hamiltonians and Gibbs sampling: An interesting future research direction is to understand the functional dependence of the quasi-quantum Gibbs Hamiltonian on the underlying stochastic dynamics. While Theorem 3.1 gives us the general shape of the interaction graph of $H_{G}$, it does not directly provide the actual form of $H_{k}$. Understanding this dependence, we might be able to engineer the local dynamics to converge to a specific, strictly local, Gibbs Hamiltonian (or an approximation of it). This would give rise to a natural family of quantum Gibbs preparation algorithms, which can be implemented on a quantum computer. Canceling high order terms in $H_{G}$ will probably require adding less local terms to stochastic dynamics. We conjecture that using quasi-local dynamics, one can reach a steady state that is a good approximation to a Gibbs state of strictly local Hamiltonian. This picture agrees with the recent Gibbs state preparation algorithms, that simulate a Lindbladian evolution with quasi-local generators [33, 24].
    Additionally, if a large family of local quantum maps give rise to a manifold of Gibbs states, then one might use a variational approach to find the local map that yields a particular Gibbs state. This approach might be particularly suitable for NISQ-type algorithms on noisy hardware, since some of the noise might be mitigated by the variational procedure [22, 28, 30].
Our second main result is given in Theorem 3.3, where we computed the expectation values for local observables within the steady state. For this purpose, instead of expanding the steady state in $\epsilon$ (Schrodinger picture), we used the expansion of the observable (Heisenberg picture). We showed that, just as for the Gibbs Hamiltonian of the steady state, the expansion of the observable in $\epsilon$ has a support that is growing with the order. Moreover, in such case, we were able to give an upper-bound for the norm of each term in the expansion. In other words, we found a positive radius of convergence for the power series of $A(\epsilon)=\sum_{k}A_{k}\epsilon^{k}$, which is an interesting result by itself [42]. As a corollary, we estimated the complexity of computing such expectation value  within the radius of convergence, which is summarized in Corollary 3.4.
Following, are two remarks on the theorem:
  1. 1.
    Strengthening the results: The assumption in Theorem 3.3 on the form of the dissipators $\mathcal{D}_{i}$ was taken for the sake of simplifying the result. This assumption can be relaxed to any ergodic dissipators which are diagonalizable, and for which the eigenoperators do not have large overlaps. Under such assumptions, the achieved bound will depend on the gap of the dissipators and the overlap between the eigenoperators. Moreover, we believe that the upper-bound on ${\|A_{k}\|}$ can be strengthened by employing different techniques other than the $Q_{1}$ norm in the Wauli basis, which introduces an undesirable exponential dependence on the support of the observable $A$. Such strengthening will turn Theorem 3.5 into a genuine ‘decay of correlations’ statement.
  2. 2.
    Complexity thresholds: By the results of Ref. [11], in the $\epsilon=1$ limit, the steady state can be the ground state of any local frustration-free Hamiltonian, and therefore estimating expectation values there up to an error of $1/\operatorname{poly}(n)$ is a $\mathrm{QMA}_{1}$-hard problem. On the other hand, by Theorem 3.3 and Corollary 3.4, for $\epsilon<\epsilon_{0}=1/e^{6}$, this problem is almost in P as it is quasi-polynomial, and therefore not expected to be QMA-hard (or even NP-hard). This implies the existence of a complexity crossover threshold somewhere in the range $[\epsilon_{0},1]$ and raises several questions: where exactly does this threshold lie? in particular, is it strictly smaller than $1$? if the answer is yes, then this might imply an interesting type of noise-resilience for such systems. It is also natural to ask if (and how) the threshold depends on underlying graph $G$. The fact that the bound $\epsilon_{0}=1/e^{6}$ in Theorem 3.3 is independent of $G$ is somewhat surprising, if we compare it other bounds of the same flavour, such as the Lieb-Robinson bound [43, 44]. Another interesting question is whether there exists a BQP-hard threshold between $\epsilon_{0}$ and the $\textrm{QMA}_{1}$ threshold? A related question is how the mixing time of the system depends on $\epsilon$. Is there an $\tilde{\epsilon}\in(\epsilon_{0},1)$ such that for $\epsilon<\tilde{\epsilon}$ the mixing time is polynomial in the system size, and therefore the problem of approximating local expectations in the steady state is in BQP? Finally, it would be interesting to understand how do all these different complexity regimes relate to the structural properties of $\rho_{\infty}$, e.g., correlation length, entanglement, efficient representability by a tensor-network, etc?
  3. 3.
    Finding a better algorithm for approximating ${\langle{A}\rangle}$: A better algorithm than in Corollary 3.4 might be devised, reducing the quasi-polynomial time complexity to a polynomial time complexity, as we explain in the following. The quasi-polynomial time complexity in Corollary 3.4 is caused by the growth in support of $A_{k}$, which is contained in the $k$-radius ball around $\operatorname{supp}(A)$. Since we consider graphs that sit on a $D$-dimensional lattice, the size of such ball is proportional to $O(k^{D})$, which leads to the $1/\delta$ quasi-polynomial running time $e^{O(k^{D})}$ when $k=\log(1/\delta)$. However, it can be seen from the proof of the theorem that $A_{k}$ is in fact a sum over $O(k)$-local Pauli (or Wauli) strings. Instead of summing all these strings, it might be possible to devise some sort of an importance-sampling algorithm for these strings (e.g., using ideas from Ref. [45]), thereby approximating the expectation value within the steady state. We leave this for future research.
7 Acknowledgments
We thank Yosi Avron, Yariv Kafri, Howard Nuer and Dorit Aharonov for fruitful discussions.
This research project was supported by the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant. It is also supported by the Israel Science Foundation (ISF) under the Individual Research Grant No. 1778/17 of I.A. and the joint Israel-Singapore NRF-ISF Research Grant No. 3528/20. We acknowledge the support of the Helen Diller Quantum Center.
Appendix A Proof of Lemma 4.1
To prove Lemma 4.1, we use the following fact on quantum channels:
Fact A.1 (Ergodicity - proposition 3 of Ref. [37])
A quantum channel $\mathcal{E}$ given in the Kraus form $\mathcal{E}(\rho)=\sum_{k=1}^{m}F_{k}\rho F_{k}^{\dagger}$ has a unique, full rank steady state $\sigma$ and no other eigenvalues for which $|\lambda|=1$ if and only if there exists $k\in\mathbb{N}$ such that the following vector space of operators
$S_{k}(\mathcal{E})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname*{Span}\{F_{i_{k}}\cdot\ldots\cdot F_{i_{1}}\;;\;i_{1},\ldots,i_{k}\in\{1,\ldots,m\}\}$ (51)
spans the entire Hilbert space of operators $L(\mathcal{H})$.
To prove Bullet 1, we first consider the case when $\epsilon_{e}=0$ for any $e\in E$, i.e., ${\underline{\epsilon}}={\underline{0}}$. The channel is then of the form
$\mathcal{E}_{{\underline{0}}}(\rho)=\sum_{e=(i,j)\in E}\!\!\!\frac{p_{e}}{2}[\mathcal{D}_{i}(\rho)+\mathcal{D}_{j}(\rho)].$
Let $\mathcal{D}_{i}(\rho)=\sum_{k=1}^{m_{i}}D_{k}^{(i)}\rho D_{k}^{(i)\dagger}$ be the Kraus decomposition of $\mathcal{D}_{i}$. The Kraus operators of the full channel $\mathcal{E}_{\underline{0}}$ is the set (up to proportionality constant of each operator)
$\bigcup_{i\in V}\{D_{k}^{(i)}\;;\;k=1,\dots,m_{i}\}.$
It is easy to see that the steady state of $\mathcal{E}_{{\underline{0}}}$ is unique, and given by the tensor product of the local steady states of each local channel $\mathcal{D}_{i}$. Moreover, it is full rank, because it is tensor product of full-rank operators, and $\mathcal{E}_{\underline{0}}$ has no other eigenvalues with absolute value $1$. By Fact A.1, there is a $k\in\mathbb{N}$ such that $S_{k}(\mathcal{E}_{\underline{0}})=L(\mathcal{H})$. From the functional form of the channel in Eq. (9), it is clear that for any ${\underline{\epsilon}}\in[0,1)^{n}$, the set of Kraus operators of $\mathcal{E}_{\underline{\epsilon}}$ includes the set $\bigcup_{i\in V}\{D_{k}^{(i)}\;;\;k=1,\dots,m_{i}\}$, again up to proportionality constants, and therefore,
$S_{k}(\mathcal{E}_{\underline{0}})\subseteq S_{k}(\mathcal{E}_{\underline{\epsilon}}).$
But as $S_{k}(\mathcal{E}_{\underline{0}})=L(\mathcal{H})$, we conclude that also $S_{k}(\mathcal{E}_{\underline{\epsilon}})=L(\mathcal{H})$, and using Fact A.1, uniqueness and full-rank of $\rho_{\infty}({\underline{\epsilon}})$ is guaranteed.
To prove the second bullet, note that the steady state satisfies the following linear equation,
$\big{(}\mathcal{E}_{\underline{\epsilon}}-\mathcal{I}\big{)}(\rho)=0.$ (52)
Using standard basis vectorization, we map $\rho_{\infty}({\underline{\epsilon}})=\sum_{ij}\rho_{ij}{|{i}\rangle\langle{j}|}\to\vec{v}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{i,j}\rho_{ij}({\underline{\epsilon}}){|{i}\rangle}{|{j}\rangle}$, where we dropped the explicit dependence of $\vec{v}$ on ${\underline{\epsilon}}$ in order to keep the notation clean. Similarly, we map the channel $\mathcal{E}_{\underline{\epsilon}}$ to a square matrix $A_{\underline{\epsilon}}$, which takes Eq. (52) to the following linear matrix equation
$(A_{\underline{\epsilon}}-\mathbbm{1})\vec{v}=0.$
Note that the matrix coefficients of $A_{\underline{\epsilon}}$ are affine in $\epsilon_{1},\ldots,\epsilon_{n}$. To solve this linear equation, we use standard Gaussian elimination, which involves subtracting from each row a multiple of another row, to reach a row canonical form [46]. Since by Bullet 1, for any ${\underline{\epsilon}}\in[0,1)^{n}$ there is a unique steady state, we conclude that the co-rank of $A-\mathbbm{1}$ must be $1$ (i.e., the rank of its null space is $1$), and therefore, its canonical form must look like
$(A-\mathbbm{1})\mapsto B=\left[\begin{array}[]{ccccc|ccccc}1&0&0\ldots&0&f_{1}&0&0&\ldots&0\\ 0&1&0\ldots&0&f_{2}&0&0&&0\\ &&\ddots&&&&&\ddots&\\ 0&0&0\ldots&1&f_{m}&0&0&\ldots&0\\ \hline\cr 0&0&0\ldots&0&0&1&0&\ldots&0\\ 0&0&0\ldots&0&0&0&1&&0\\ &&\vdots&&&&&\ddots&\\ 0&0&0\ldots&0&0&0&\ldots&&1\\ 0&0&0\ldots&0&0&0&\ldots&&0\end{array}\right]$
Since Gaussian elimination involves subtracting multiple of rows and dividing each row by the first non-zero coefficient, we conclude that $f_{1},\ldots,f_{m}$ are in-fact polynomial quotient
$f_{i}(\epsilon_{1},\ldots,\epsilon_{n})=\frac{h_{i}(\epsilon_{1},\ldots,\epsilon_{n})}{g_{i}(\epsilon_{1},\ldots,\epsilon_{n})}$
with $h_{i},g_{i}$ being polynomials. As the system of equations $(A-\mathbbm{1})\vec{v}=0$ is equivalent to $B\vec{v}=0$, we conclude that
$v_{n}$ $=v_{n-1}=\dots=v_{m+2}=0,$
$v_{m+1}$ $=c,$
$v_{i}$ $=-c/f_{i}\quad\forall i=1,\ldots,m,$
where $c$ is a constant that is fixed later by the normalization condition $\mathrm{Tr}(\rho_{\infty})=1$. Note that $c$ is by itself a polynomial quotient, thus also $v_{i}$ is, namely, $v_{i}=p_{i}/q_{i}$ for some polynomials $p_{i},q_{i}$ for each $i=1,\ldots,m$. Since the coefficients of $\vec{v}$ are the entries of $\rho_{\infty}({\underline{\epsilon}})$, and since we know that $\rho_{\infty}({\underline{\epsilon}})$ exists for any ${\underline{\epsilon}}\in[0,1)^{n}$, we are guaranteed that there are no poles for any $q_{i}$ in the open cube $[0,1)^{n}$, and therefore the entries of $\rho_{\infty}({\underline{\epsilon}})$ are analytical in an open set containing $[0,1)^{n}$ [47].
Appendix B $Q_{1}$-norm and $\mathcal{T}$ properties
Here we provide the proofs for properties of the $Q_{1}$-norm and the super-operator $\mathcal{T}$ that were stated and used in the proof of Theorem 3.3
  • Proof of Lemma 4.6:
    1. 1.
      Recalling that in the dual Wauli basis, $\tilde{Q}_{0}=\mathbbm{1}$, the proof follows immediately.
    2. 2.
      As $O\in S_{k}$, its dual Wauli expansion is $O=\sum_{|{\underline{\smash{\alpha}}}|\leq k}c_{\underline{\smash{\alpha}}}\tilde{Q}_{\underline{\smash{\alpha}}}$. Therefore,
      ${\|O\|}_{\infty}=\Big{\|}\sum_{|{\underline{\smash{\alpha}}}|\leq k}c_{\underline{\smash{\alpha}}}\tilde{Q}_{\underline{\smash{\alpha}}}\Big{\|}_{\infty}\leq\sum_{|{\underline{\smash{\alpha}}}|\leq k}|c_{\underline{\smash{\alpha}}}|\cdot{\|\tilde{Q}_{\underline{\smash{\alpha}}}\|}_{\infty}\leq(1+\lambda)^{k}\cdot\|O\|_{Q_{1}},$
      where in the last step we used the fact that for a single qubit dual-Wauli element $\tilde{Q}_{\alpha}$, the highest possible eigenvalue (obtained for $\alpha=3$) is $1+\lambda$, and therefore the highest eigenvalue of $\tilde{Q}_{\underline{\smash{\alpha}}}$ which is supported on at most $k$ qubits is bounded by $(1+\lambda)^{k}$.
    3. 3.
      For inequality (23), we expand $O=\sum_{\underline{\smash{\alpha}}}c_{\underline{\smash{\alpha}}}\tilde{Q}_{\underline{\smash{\alpha}}}$ and use the Hilbert-Schmidt product to express $c_{\underline{\smash{\alpha}}}=\mathrm{Tr}(Q_{\underline{\smash{\alpha}}}\cdot O)$. Then
      $\|O\|_{Q_{1}}$ $=\sum_{{\underline{\smash{\alpha}}}}|c_{\alpha}|=\sum_{\underline{\smash{\alpha}}}|\mathrm{Tr}(Q_{\underline{\smash{\alpha}}}\cdot O)|\leq\sum_{\underline{\smash{\alpha}}}{\|Q_{\underline{\smash{\alpha}}}\|}_{1}\cdot{\|O\|}_{\infty}\leq 4^{\ell}\cdot{\|O\|}_{\infty},$
      where in the first inequality we used Holder’s inequality, and in the last inequality we used the fact that the support of $O$ is $\ell$, and therefore its expansion contains at most $4^{\ell}$ terms. In addition, for every ${\underline{\smash{\alpha}}}$, it is easy to see that ${\|Q_{\underline{\smash{\alpha}}}\|}_{1}=1$.
To prove Lemma 4.7 we need the following fact:
Fact B.1
For a CPTP map $\mathcal{E}$ and an hermitian operator $O$ we have that
${\|\mathcal{E}^{*}(O)\|}_{\infty}\leq{\|O\|}_{\infty}.$
  • Proof of fact B.1:
    ${\|\mathcal{E}^{*}(O)\|}_{\infty}=\max_{\rho\in D(\mathcal{H})}|\mathrm{Tr}(\mathcal{E}^{*}(O)\rho)|=\max_{\rho\in D(\mathcal{H})}|\mathrm{Tr}(O\mathcal{E}(\rho))|\leq\max_{\rho^{\prime}\in D(\mathcal{H})}|\mathrm{Tr}(O\rho^{\prime}))|={\|O\|}_{\infty},$
    where $D(\mathcal{H})$ denotes the set of quantum states, and in the first and last equalities we used the fact that for a hermitian operator $A$, ${\|A\|}_{\infty}=\max_{\rho\in D(\mathcal{H})}|\mathrm{Tr}(A\cdot\rho)|$.
  • Proof of lemma 4.7:
    Expand $O=\sum_{\underline{\smash{\alpha}}}c_{\underline{\smash{\alpha}}}\tilde{Q}_{\underline{\smash{\alpha}}}$ and decompose each string ${\underline{\smash{\alpha}}}$ to ${\underline{\smash{\alpha}}}_{a},{\underline{\smash{\alpha}}}_{b}$, where ${\underline{\smash{\alpha}}}_{a}$ acts on spins in the support of $\mathcal{E}$ and ${\underline{\smash{\alpha}}}_{b}$ on the rest of the system. Then for every ${\underline{\smash{\alpha}}}$, decompose $\tilde{Q}_{\underline{\smash{\alpha}}}=\tilde{Q}_{{\underline{\smash{\alpha}}}_{a}}\otimes\tilde{Q}_{{\underline{\smash{\alpha}}}_{b}}$, and upper-bound
[…]
    Finally, using the fact that $\lambda\leq 1$, we conclude that $4^{k}(1+\lambda)^{k}\leq 2^{3k}$.
  • Proof of Claim 4.8:
    We start by deriving an explicit formula for the application of $\mathcal{T}$ on a dual Wauli operator $\tilde{Q}_{\underline{\smash{\alpha}}}$. Applying $\mathcal{E}_{1}^{*}$ on Eq. (39) we get:
    $\mathcal{T}(\tilde{Q}_{\underline{0}})$ $=\mathcal{E}_{1}^{*}(\tilde{Q}_{\underline{0}})=0,\qquad$ (53)
    $\mathcal{T}(\tilde{Q}_{\underline{\smash{\alpha}}})$ $=\frac{1}{q_{\underline{\smash{\alpha}}}}\mathcal{E}_{1}^{*}(\tilde{Q}_{\underline{\smash{\alpha}}})=\frac{1}{q_{\underline{\smash{\alpha}}}}\sum_{e\in E}p_{e}\mathcal{F}^{*}_{e}(\tilde{Q}_{\underline{\smash{\alpha}}})-\frac{1-q_{\underline{\smash{\alpha}}}}{q_{\underline{\smash{\alpha}}}}\tilde{Q}_{\underline{\smash{\alpha}}},\qquad({\underline{\smash{\alpha}}}\neq{\underline{0}}).$ (54)
    where in the first equation we have used the fact that dual quantum channels are unital, therefore $\mathcal{E}_{1}^{*}$, being a difference of such channels, maps the identity $\tilde{Q}_{\underline{0}}$ to 00. Similarly, $\mathcal{F}_{e}^{*}$ is unital and since it is $2$-local (acting on an edge), we must have $|\operatorname{supp}(\mathcal{F}^{*}_{e}(\tilde{Q}_{\underline{\smash{\alpha}}}))|\leq 1+|\operatorname{supp}(\tilde{Q}_{\underline{\smash{\alpha}}})|$. As the other terms of $\mathcal{T}$ cannot increase the support, we get that for any $O\in\mathcal{S}_{k}$, by expanding it in the dual-Waulis, we must have $\mathcal{T}(O)\in\mathcal{S}_{k+1}$ by the linearity of the channel. Moreover, given that $O$ is supported in a subset $A$, or alternatively, any dual-Wauli $\tilde{Q}_{\alpha}$ in its expansion is supported in $A$, under the same reasoning we would get that $\operatorname{supp}(\tilde{Q}_{\alpha})$ is inside $A\cup e$ (provided that $e$ intersects $A$). This concludes Bullet 1.
    Next, let $E_{0}$ denote the subset of edges $e=(i,j)$ for which $\alpha_{i}=\alpha_{j}=0$, i.e., the edges that are fully outside the support of ${\underline{\smash{\alpha}}}$, and let $E_{1}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}E\setminus E_{0}$. For $e\in E_{0}$, we have $\mathcal{F}_{e}^{*}(\tilde{Q}_{\underline{\smash{\alpha}}})=\tilde{Q}_{\underline{\smash{\alpha}}}$ since $\tilde{Q}_{\underline{\smash{\alpha}}}$ reduces to the identity on qubits $i,j$. Thus, breaking the sum to $E_{0}$ and $E_{1}$ gives
    $\sum_{e\in E}p_{e}\mathcal{F}^{*}_{e}(\tilde{Q}_{\underline{\smash{\alpha}}})=\big{(}1-p_{E_{1}}\big{)}\tilde{Q}_{\underline{\smash{\alpha}}}+\sum_{e\in E_{1}}p_{e}\mathcal{F}^{*}_{e}(\tilde{Q}_{\underline{\smash{\alpha}}}),$
    where we use the notation $p_{E_{1}}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\sum_{e\in E_{1}}p_{e}$ and the fact that $\{p_{e}\}$ form a probability distribution. Plugging this into Eq. (54) for ${\underline{\smash{\alpha}}}\neq{\underline{0}}$ get
    $\mathcal{T}(\tilde{Q}_{\underline{\smash{\alpha}}})$ $=\frac{1}{q_{\underline{\smash{\alpha}}}}\sum_{e\in E_{1}}p_{e}\mathcal{F}^{*}_{e}(\tilde{Q}_{\underline{\smash{\alpha}}})+\left(1-\frac{p_{E_{1}}}{q_{\underline{\smash{\alpha}}}}\right)\tilde{Q}_{\underline{\smash{\alpha}}}.$ (55)
    With this expression in hand, we proceed to prove Bullet 2. Let $O$ be a general observable expanded in the dual-Waulis $O=\sum_{\underline{\smash{\alpha}}}c_{\underline{\smash{\alpha}}}\tilde{Q}_{\underline{\smash{\alpha}}}$. Before bounding $\|\mathcal{T}(O)\|_{Q_{1}}$, we note following:
    1. i
      For any dual Wauli $\tilde{Q}_{\underline{\smash{\alpha}}}$, clearly $\|\tilde{Q}_{\underline{\smash{\alpha}}}\|_{Q_{1}}=1$
    2. ii
      Therefore, by Lemma 4.7, $\|\mathcal{F}^{*}_{e}(\tilde{Q}_{\underline{\smash{\alpha}}})\|_{Q_{1}}\leq 2^{6}\|\tilde{Q}_{\underline{\smash{\alpha}}}\|_{Q_{1}}=64$ as $\mathcal{F}^{*}_{e}$, is $2$-local.
    3. iii
      Given the relation between $q_{i}$ and $p_{e}$ from Eq. (25), it follows that $p_{E_{1}}\leq 2q_{{\underline{\smash{\alpha}}}}$, since:
      $p_{E_{1}}=\sum_{e\in E_{1}}p_{e}\leq\sum_{i\in\operatorname{supp}({\underline{\smash{\alpha}}})}\sum_{e\ni i}p_{e}=2\sum_{i\in\operatorname{supp}({\underline{\smash{\alpha}}})}q_{i}=2q_{{\underline{\smash{\alpha}}}}$
    Using these, together with (53) and (55), we obtain the desired bound:
    ${\|\mathcal{T}(O)\|}_{Q_{1}}$ $\leq\sum_{{\underline{\smash{\alpha}}}}|c_{\underline{\smash{\alpha}}}|\cdot{\|\mathcal{T}(\tilde{Q}_{\underline{\smash{\alpha}}})\|}_{Q_{1}}$
    $\leq\sum_{{\underline{\smash{\alpha}}}\neq{\underline{0}}}|c_{\underline{\smash{\alpha}}}|\cdot\left(\frac{1}{q_{\underline{\smash{\alpha}}}}\sum_{e\in E_{1}}p_{e}{\|\mathcal{F}^{*}_{e}(\tilde{Q}_{\underline{\smash{\alpha}}})\|}_{Q_{1}}+\left|1-\frac{p_{E_{1}}}{q_{\underline{\smash{\alpha}}}}\right|\cdot{\|\tilde{Q}_{\underline{\smash{\alpha}}}\|}_{Q_{1}}\right)$
    $\leq\sum_{{\underline{\smash{\alpha}}}\neq{\underline{0}}}|c_{\underline{\smash{\alpha}}}|\left(\frac{p_{E_{1}}}{q_{\underline{\smash{\alpha}}}}\cdot 64+\left|1-\frac{p_{E_{1}}}{q_{\underline{\smash{\alpha}}}}\right|\right)$
    $\leq 130\cdot\sum_{{\underline{\smash{\alpha}}}}|c_{\underline{\smash{\alpha}}}|=130\cdot\|O\|_{Q_{1}}.$